Day shift summary, 12/11/99

AG (ag@fnal.gov)
Sat, 11 Dec 1999 16:32:18 -0500 (EST)

On shift: AG

0700 COSMICS run 15197 is still in progress.

1250 ktev3 processes fail to run on xktev6c terminal
-> reboot it as per Theo's suggestion

1430 Ed phones and asks to stop the run, so that he can download constants
for channel 1096 from two days earlier, and update "PMT gain drift"
plot.

1500 Updated csq6 and cspd15198. Ed is finished, so start another COSMICS
run. He also mentioned that it would be necessary to monitor
channels:

799 1096 2847 2091

when beam comes back.

1510 Did drt fresh start to restore mon_disp, which died in the previous
run.